# 3DGS.zip: A survey on 3D Gaussian Splatting Compression Methods

### v4

@@ 4.1.2 Structured Representation of Gaussians
One approach to structure 3D Gaussians of a scene is the Anchor-Based Representation as introduced in Scaffold-GS [29]. These anchors act as representative points for clusters of Gaussians in close proximity. Instead of storing attributes for each individual Gaussian, the attributes are associated with the anchors. A neural network then predicts the attributes of the Gaussians within each cluster based on the anchor’s attributes. This approach effectively reduces the overall number of parameters required to represent the scene. In HAC [8], the Anchor-Based Representation is enhanced through the inclusion of a Hash-Grid Assisted Context. The primary concept involves simultaneously learning a structured, compact hash grid for context modeling of anchor attributes. For each anchor, the anchor’s position is used to query the hash grid and retrieve an interpolated hash feature, which subsequently predicts the value distributions of the anchor attributes to aid in entropy coding for a highly efficient representation.
@@
An alternative structuring methodology is employed by SOG [35] and also used in gsplat[19], where unstructured Gaussians are mapped onto a structured 2D grid to spatially organize Gaussians with similar attributes closely, enhancing the smoothness of attribute values. This configuration allows for efficient compression using standard image compression techniques. To further improve smoothness and compressibility, a smoothing regularization term can be added during training to promote locally smooth configurations of Gaussians on the grid.
The authors in CompGS [37] suggest an approach for organizing the unstructured 3D Gaussians by sorting them according to one of the quantized indices and subsequently storing them employing Run-Length Encoding (RLE).
IGS[45] proposes a hybrid approach that integrates explicit point clouds with implicit feature embeddings, focusing on the encoding of Gaussian attributes within a multi-resolution tri-plane architecture. This structure utilizes 2D feature grids at varying resolutions to represent scene space, enhancing spatial correlations and allowing a continuous representation of the spatial domain. A progressive training strategy is adopted, starting with coarse predictions at lower levels and refining them through higher-level tri-planes, leading to a compact and efficient 3D scene representation.
@@ 4.1.3 Quantitative Comparison of Compression Methods
This section presents the comparative analysis of 3DGS compression methods across four datasets: Tanks and Temples, Mip-NeRF 360, Deep Blendingand, Synthetic NeRF. More details on the datasets are provided in Sec. 5.1. As shown in Table 1, the metrics used to evaluate the performance (see Sec. 5.2) of each compression method are PSNR, SSIM, LPIPS and model size measured in megabytes. We calculate an additional rank that represents the average rankings of the methods across all available datasets. To determine the compression dataset ranks $rank_{s}$, the ranks for the quality metrics PSNR, SSIM, and LPIPS are equally weighted with the model size. Thus, each quality metric contributes one-sixth to the ranking, while the model size accounts for the remaining half:
@@
The methods evaluated show significant variation in file size, with some achieving high compression rates at the cost of visual quality, while others strike a balance between compression and maintaining higher fidelity. While our proposed ranking puts HAC-highrate on the first place, high variations between datasets and quality metrics show that there is not one winning compression strategy. Depending on the application and goals, different approaches should be considered. If file size is the main concern HAC-lowrate, IGS low, and SOG w/o SH are the best-performing compression methods, achieving file sizes under 10 MB with acceptable levels of visual quality. HAC-lowrate [8] is particularly effective, offering aggressive compression (8.5 MB) with minimal loss in quality. SOG w/o SH [35] achieves similar compression (9.3 MB) but with slightly lower fidelity. HAC-highrate [8] provides an excellent balance between compression and quality, maintaining a relatively small size with high PSNR, making it a versatile option. IGS high [45] offers higher PSNR on the Synthetic NeRF and Deep Blending datasets but require larger file sizes, suitable for cases prioritizing quality over compression.
Figure 7 visually illustrates the trade-offs between model size and PSNR for various 3DGS compression methods across different datasets. Some approaches have additional data points, which for clarity were not included in the table. The figure illustrates the trade-offs between model size and PSNR and highlights how smaller file sizes typically result in lower PSNR, showing the balance between compression efficiency and visual quality. Further figures for the comparison of SSIM and LPIPS as well as the corresponding compaction figures can be found in Appendix B.
@@ 4.2.1 Densification
Lastly, Markov Chain Monte Carlo (MCMC) [19] uses a completely different approach for densifying the 3DGS scene. Instead of cloning and splitting the individual splats, they sample a fixed number of Gaussians from a learned probability distribution. Specifically, they add a noise term to the position of the Gaussians to incorporate exploration during the training, while they also re-spawn Gaussians that drop below a opacity threshold at randomly at Gaussians with high opacity. This way, the method improves the quality of regions with few initial Gaussian splats and the number of Gaussians can directly be controlled.
@@ 6 Discussion
The field of 3D Gaussian Splatting (3DGS) has seen rapid advancements, with significant improvements in compression and compaction techniques aimed at overcoming the high memory demands traditionally associated with 3DGS. This state-of-the-art report reveals key insights into these developments, providing a foundation for future optimization and application of 3DGS in real-time rendering.
@@
A crucial finding is that a good compaction strategy can be an improvement to most methods. Adaptive Density Control, with densification and pruning can significantly improve visual quality while reducing the memory footprint. This is vital for applications in resource-constrained environments, such as mobile devices or VR headsets, where memory and processing power are limited but visual quality matters to the user.
Structured representations play a key role in achieving good compression ratios. Techniques like Anchor-Based Representations and Hash-Grid Assisted Contexts organize Gaussian data to reduce redundancy and improve compressibility. This approach keeps 3DGS models compact and maintains high visual quality, vital for efficiently storing and retrieving large, complex scenes. Furthermore, strategies like vector quantization and basic attribute pruning show that simplicity, when combined with a focus on key optimization strategies, can yield highly efficient models without sacrificing too much visual fidelity.
@@
Finally, the trade-off between compression efficiency and visual quality remains a central challenge. Approaches like HAC-highrate excel in achieving high-quality compression but at the cost of higher computational overhead. On the other hand, more aggressive compression techniques such as SOG w/o SH reduce memory usage significantly but may result in noticeable quality degradation, especially in scenes requiring fine detail. This trade-off highlights the need for flexible solutions that can be tuned based on the specific requirements of the application, whether it’s focused on minimizing storage or maximizing visual realism.
@@ 7 Conclusion and Future Directions
This state-of-the-art report reveals the importance of continuing to refine and standardize 3D Gaussian Splatting (3DGS) compression and compaction techniques to make them more accessible and widely applicable. Current methods provide effective solutions for reducing the memory and computational load of 3DGS, but there is still room for improvement in terms of scalability, ease of use, and adaptability across different platforms.
One of the most promising future directions is extending 3DGS to dynamic scenes, where objects and lighting conditions change over time. This advancement would enable real-time simulations and enhance interactive applications such as gaming and virtual reality. Another improvement could be the creation of multi resolution models with support for level-of-detail (LOD) scaling. Such models would optimize performance by allowing different parts of a scene to be rendered with varying levels of detail, depending on real-time requirements. Quantization-aware training and the development of shared codebooks across scenes or applications could further improve compression efficiency, allowing for reduced redundancy and more effective memory usage.
Together, these advancements will help expand the applicability of 3DGS, making it a more versatile and efficient tool for addressing future challenges in computational graphics.
@@ 8.1 HAC: Hash-grid Assisted Context for 3D Gaussian Splatting Compression (HAC)
8.3 Implicit Gaussian Splatting with Efficient Multi-Level Tri-Plane Representation (IGS)
This method[45] introduces a hybrid representation for splatting-based radiance fields, where Gaussian primitives are separated into explicit point cloud and implicit attribute features. The attribute features are encoded using a multi-resolution, multi-level tri-plane architecture integrated with a residual-based rendering pipeline. It employs a level-based progressive training scheme for joint optimization of point clouds and tri-planes, starting with coarse attributes and refining them with higher-level details. Spatial regularization and a bootstrapping scheme are applied to enhance the consistency and stability of the Gaussian attributes during training.
8.4 Compact3D: Compressing Gaussian Splat Radiance Field Models with Vector Quantization (CompGS)
@@ 8.9 Compressed 3D Gaussian Splatting for Accelerated Novel View Synthesis (Compressed3D)
The authors propose a compressed 3D Gaussian splat representation [38] consisting of three main steps: 1. sensitivity-aware clustering, where scene parameters are measured according to their contribution to the training images and encoded into compact codebooks via sensitivity-aware vector quantization; 2. quantization-aware fine-tuning, which recovers lost information by fine-tuning parameters at reduced bit-rates using quantization-aware training; and 3. entropy encoding, which exploits spatial coherence through entropy and run-length encoding by linearizing 3D Gaussians along a space-filling curve. Furthermore, a renderer for the compressed scenes utilizing GPU-based sorting and rasterization is proposed, enabling real-time novel view synthesis on low-end devices.
@@ 9.1 Mini-Splatting: Representing Scenes with a Constrained Number of Gaussians (Mini-Splatting)
Mini-Splatting [12] enhances Gaussian distribution through Blur Split, which refines Gaussians in blurred regions, and Depth Reinitialization, which repositions Gaussians based on newly generated depth points, calculated from the mid-point of ray intersections with Gaussian ellipsoids, thus avoiding artefacts from alpha blending. For simplification, Intersection Preserving retains Gaussians with the greatest visual impact, while Sampling maintains geometric integrity and rendering quality, reducing complexity.

### v5

@@ 4.1.2 Structured Representation of Gaussians
One approach for structuring 3D Gaussians whithin a scene is the Anchor-Based representation, as introduced in Scaffold-GS [34]. From the initial Structure from Motion (SfM) derived point cloud a scene is first voxelized and each voxel center is then treated as an anchor. Each anchor is associated with a local context and learnable offsets, acting as representative point which can spawn new Gaussians in close proximity. Subsequently a neural network predicts the attributes i.e. opacity, color and covariance, of the associated Gaussians based on anchor features and viewing conditions. This approach effectively reduces the overall number of parameters required to represent the scene. ContextGS [50] builds upon the Anchor representation introduced in Scaffold-GS [34] but divides the anchor points into hierarchical levels, from coarse to fine. The anchors are encoded progressively, starting from the coarsest level. The decoded values of the anchors at a coarser level are then used to predict the distribution of nearby anchors at the next finer level, using an MLP. The hierarchical anchor approach efficiently exploits the spatial relationships between anchors. In HAC [9], the Anchor-Based representation is enhanced through the inclusion of a Hash-Grid Assisted Context. The primary concept involves simultaneously learning a structured, compact hash grid for context modeling of anchor attributes. For each anchor, the anchor’s position is used to query the hash grid and retrieve an interpolated hash feature, which subsequently predicts the value distributions of the anchor attributes to aid in entropy coding for a highly efficient representation. While also based on an Anchor-Based representation with anchor primitives and coupled primitives the primitives differ in their structure CompGS [32]. Anchor primitives serve as reference and contain geometry attributes (location and covariance) and reference embeddings. Coupled primitives only contain residual embeddings to capture deviations. The attributes of coupled primitives are predicted by warping corresponding anchor primitives using affine transforms derived from anchor and coupled primitive embeddings.
@@
An alternative structuring methodology is employed by SOG [40] and also used in gsplat[22], where unstructured Gaussians are mapped onto a structured 2D grid to spatially organize Gaussians with similar attributes closely, enhancing the smoothness of attribute values. This configuration allows for efficient compression using standard image compression techniques. To further improve smoothness and compressibility, a smoothing regularization term can be added during training to promote locally smooth configurations of Gaussians on the grid. An other approach CodecGS [27] applies feature planes for efficient attribute representation and uses a standard video coding technique for feature plane compression. The authors propose a tri-plane architecture to predict Gaussian attributes instead of storing them directly. A progressive training strategy allows to capture coarse geometric information initially and gradually add finer details. The feature planes are optimized by applying a block-wise discrete cosine transform (DCT) allowing the model to leverage the spatial correlations within the feature planes as conventional image signals.
The authors in Compact3D [42] suggest an approach for organizing the unstructured 3D Gaussians by sorting them according to one of the quantized indices and subsequently storing them employing Run-Length Encoding (RLE).
@@ 4.1.3 Quantitative Comparison of Compression Methods
This section presents the comparative analysis of 3DGS compression methods across four datasets: Tanks and Temples, Mip-NeRF 360, Deep Blending and, Synthetic NeRF. More details on the datasets are provided in Sec. 5.1. For better comparison Table 1 only includes approaches that mainly focus on compression, we provide a separate Table 2 to compare approaches that mainly focus on compaction (i.e. densification and pruning). Furthermore some approaches were missing the necessary data for quantitative comparison, approaches that we consider worth mentioning nevertheless are included in the survey but not in the tables. As shown in Table 1, the metrics used to evaluate the performance (see Sec. 5.2) of each compression method are PSNR, SSIM, LPIPS and model size measured in megabytes.
The objectives for compressing 3DGS vary depending on the application, with some requiring minimal model size, while others prioritize a smaller size alongside optimal perceptual quality. As there is no definitive winner that excels across all categories, we introduce a simple rank, reflecting the average rankings of the methods across all available datasets, thereby offering general guidance on the overall performance of the approaches. To determine the compression dataset $rank_{s}$, the ranks for the quality metrics PSNR, SSIM, and LPIPS are equally weight with the model size. Consequently, each quality metric contributes one-sixth to the overall ranking, while the model size accounts for the remaining half:
@@
The methods evaluated show significant variation in file size, with some achieving high compression rates at the cost of visual quality, while others strike a balance between compression and maintaining higher fidelity. While our proposed ranking puts ContextGS-lowrate [50] on the first place, high variations between datasets and quality metrics show that there is not one winning compression strategy. Depending on the application and goals, different approaches should be considered. When size is the main concern CodecGS[27] stands out for its very small file size (e.g., just 7.8 MB on Tanks and Temples), all while maintaining PSNR/SSIM near the top tier. Furthermore, HAC-lowrate [9], SOG w/o SH [40] and ContextGS_lowrate [50] are the best-performing compression methods, achieving file sizes under 10 MB on the Tanks and Temples dataset with acceptable levels of visual quality.
When the emphasis is on perceptual quality (LPIPS) gsplat [55] proves to be an excellent choice, alongside ContextGS_highrate [50] and HAC_highrate [9]. It is noteworthy that the original 3DGS-30K [21] exhibits the best LPIPS values on both the Mip-NeRF 360  and the Deep Blending  dataset, at the cost of being up to 70 times larger in size.
Taken together, the top-ranking methods (ContextGS [50], HAC [9], CodecGS [27], gsplat [55]) achieve a balanced compromise among PSNR, SSIM, LPIPS, and size, with slight differences depending on whether absolute fidelity or smaller storage is the priority. The prevailing conclusion is that methods integrating context modeling (ContextGS [50]) or hierarchical approaches (HAC [9]) tend to dominate in terms of image quality, whereas CodecGS and the "lowrate" variants are particularly robust for model compression.
Figure 12 illustrates the trade-offs between model size and performance measured by PSNR for various 3DGS compression methods across different datasets. The curves highlight how smaller file sizes typically result in lower PSNR, showing the balance between compression efficiency and visual quality. Some approaches have additional data points, which for clarity were not included in the table. Further figures for the performance comparison measured by SSIM and LPIPS as well as the corresponding compaction figures can be found in Appendix B.
@@ 4.2.1 Densification
Lastly, Markov Chain Monte Carlo (MCMC) [22] uses a completely different approach for densifying the 3DGS scene. Instead of cloning and splitting the individual Gaussians, they sample a fixed number of Gaussians from a learned probability distribution. Specifically, they add a noise term to the position of the Gaussians to incorporate exploration during the training, while they also re-spawn Gaussians that drop below an opacity threshold randomly at Gaussians with high opacity. This way, the method improves the quality of regions with few initial Gaussian primitives and the number of Gaussians can directly be controlled.
@@ 5.3 Testing Protocols
Table 3 provides a comparative analysis of training and evaluation sizes. The gsplat [55] methodology was employed to train Mip-NeRF 360  scenes at three different resolutions: full resolution, 1600 pixels (longest side), and resolutions downscaled by factors of 2 or 4 (longest side $<1600$ pixels). The findings indicate that the optimal Peak Signal-to-Noise Ratio (PSNR) results are achieved when the evaluation resolution corresponds precisely to the training resolution. Any deviation, whether by increasing or decreasing the evaluation resolution relative to the training resolution, results in a slight reduction in PSNR. Furthermore, utilizing both higher training and evaluation resolutions enhances PSNR. Additionally, the overall quality is enhanced when employing 1 million Gaussians compared to 360k Gaussians, which aligns with expectations, as rendering scenes with 1 million Gaussians captures more details. Consequently, it is advised to carefully select training and evaluation sizes, acknowledging that this decision affects the final PSNR outcome as well as the comparative analysis of different compression approaches.
Table 3: PSNR comparison table for training sizes vs. evaluation sizes. For this table we used gsplat[55] to train Mip-NeRF 360 scenes either in full resolution, with a resolution of 1600 pixels or scaled down with a factor of 2 or 4 (resulting in a resolution less than 1600 pixels, as used in the implementations [4, 21]). Evaluation is performed using full resolution or 1600 pixels (longest dimension).
PSNR PSNR
(eval. res. full) (eval. res. 1600px)
##\##Gaussians $\xrightarrow{}$ 360 K 1 M 360 K 1 M
[…]
scaling factor 2/4 25.58 25.84 26.40 27.02
[…]
@@ 6 Discussion
The field of 3D Gaussian Splatting (3DGS) has seen rapid advancements, with significant improvements in compression and compaction techniques aimed at overcoming the high storage demands traditionally associated with 3DGS. While the original implementation of 3D Gaussian Splatting [21] trained on scenes from the Tanks and Templesdataset resulted in a data size exceeding 400 MB, the most effective compression methods achieve more than a 40-fold reduction in size, all the while enhancing visual quality. This state-of-the-art report reveals key insights into these developments, providing a foundation for future optimization and application of 3DGS in real-time rendering.
@@
A crucial finding is that a well-implemented compaction strategy can be an improvement to most methods. Adaptive Density Control (ADC), with densification and pruning, has proven to be effective in significantly enhancing visual quality while reducing the storage footprint. This is vital for applications in resource-constrained environments, such as mobile devices or VR headsets, where storage, memory and processing power are limited but visual fidelity matters to the user. Structured representations also play a key role in achieving efficient compression. While various compression techniques have been developed independently, their interplay remains an open research area. Understanding the interactions between different methods can lead to more effective strategies, such as vector quantization reducing redundancy by encoding similar Gaussians with shared representations, or structured representations like octrees and hash grids spatially organizing Gaussians to improve compression. Further research into integrating vector quantization within structured representations could lead to more effective encodings.
Techniques that prune Gaussian attributes, such as spherical harmonics (SH) coefficients, must be balanced against compaction strategies like ADC, which reduces the number of Gaussians. Determining the optimal trade-off between reducing attributes and minimizing Gaussian count remains an important area of study. Some compression methods prioritize extreme size reduction through entropy coding and quantization, while others focus on preserving fidelity using hierarchical encoding. Hybrid approaches that dynamically adjust compression levels based on scene complexity could be highly beneficial. Compression techniques must balance reducing file size with maintaining rendering accuracy. Neural feature encoding techniques, such as tri-planes or hash grids, significantly reduce storage needs but require additional computation during rendering. In contrast, explicit representations with precomputed attributes demand more storage but enable real-time rendering with minimal overhead. Lossy compression techniques like pruning or aggressive quantization can introduce artifacts, particularly in view-dependent rendering. Knowledge distillation techniques that encapsulate high-frequency details into lower-order representations offer a promising approach to maintaining perceptual quality while reducing complexity.
Another key challenge is the lack of standard benchmarks for evaluating 3DGS compression methods. Current studies use different datasets, metrics, and experimental conditions, making direct comparisons difficult. In this survey we propose basic guidelines to enable comparison but establishing a unified benchmarking framework with standardized datasets and evaluation protocols would enhance reproducibility and facilitate progress in the field. In addition, methods should be evaluated not only on file size and reconstruction quality but also on real-time performance and energy efficiency, particularly for deployment on mobile and embedded systems.
@@
Finally, the trade-off between compression efficiency and visual quality remains a central challenge. Approaches like HAC-highrate excel in achieving high-quality compression but at the cost of higher computational overhead. On the other hand, more aggressive compression techniques such as SOG w/o SH reduce memory usage significantly but may result in noticeable quality degradation, especially in scenes requiring fine detail. This trade-off highlights the need for flexible solutions that can be tuned based on the specific requirements of the application, whether it’s focused on minimizing storage or maximizing visual realism. As the field advances, hybrid approaches that integrate machine learning-driven optimizations may offer new opportunities. Deep learning models could be trained to predict optimal compression strategies based on scene characteristics, dynamically selecting techniques that maximize efficiency while preserving quality.
@@ 7 Conclusion and Future Directions
3D Gaussian Splatting (3DGS) has emerged as a powerful alternative to neural radiance fields, offering explicit control over scene elements while achieving high rendering fidelity. However, its practicality remains constrained by storage and computational costs. Scalability, ease of use, and adaptability across different platforms can still be improved. This state-of-the-art-report provides an overview of existing compression and compaction techniques, highlighting key advances and challenges in the field and the importance of continuing to refine and standardize 3D Gaussian Splatting (3DGS) compression and compaction techniques to make them more accessible and widely applicable.
Most current 3DGS compression techniques are designed for static scenes. However, real-world applications increasingly demand dynamic and interactive capabilities. Existing methods rely on pre-trained models, limiting adaptability. Developing real-time adaptive compression strategies that adjust based on scene complexity such as changing objects or lighting conditions over time, or changing hardware constraints would be a valuable research direction. Further advancement could enable real-time simulations and enhance interactive applications such as gaming and virtual reality.
Another improvement could be the creation of multi resolution models with support for level-of-detail (LOD) scaling. Such models would optimize performance by allowing different parts of a scene to be rendered with varying levels of detail, depending on real-time requirements. Octrees or feature grids, provide scalable solutions, but ensuring seamless transitions between different levels of detail without introducing artifacts remains a challenge. Quantization-aware training and the development of shared codebooks across scenes or applications could further improve compression efficiency, allowing for reduced redundancy and more effective memory usage.
Together, these advancements will help expand the applicability of 3DGS to encompass to real-time applications, immersive environments, or large-scale scene representations making 3DGS a more versatile and efficient tool for addressing future challenges in the field of computational graphics.
@@ 8 Short Summaries of Included Compression Approaches
8.1 ContextGS: Compact 3D Gaussian Splatting with Anchor Level Context Model (ContextGS)
The authors of ContextGS [50] proposes the first auto-regressive model at the anchor level for 3DGS compression. This work divides anchors into different levels and the anchors that are not coded yet can be predicted based on the already coded ones in all the coarser levels, leading to more accurate modeling and higher coding efficiency. To further improve the efficiency of entropy coding, a low-dimensional quantized feature is introduced as the hyperprior for each anchor, which can be effectively compressed. This work can be applied to both Scaffold-GS and vanilla 3DGS.
@@ 8.2 HAC: Hash-grid Assisted Context for 3D Gaussian Splatting Compression (HAC)
8.3 Compression of 3D Gaussian Splatting with Optimized Feature Planes and Standard Video Codecs (CodecGS)
This method [27] introduces an effective approach for compressing 3D Gaussian Splatting by employing optimized feature planes and integrating them with standard video codecs. More specifically CodecGS introduces progressive tri-planes, where the tri-plane takes 3D point positions x as input and predicts the corresponding attributes for each point. A two-phase training, starting with standard 3DGS [21] training followed by feature plane training ensures densification is consistent. Progressive training addresses the instability of feature plane training due to sparse 3DGS signals. DCT entropy modeling is employed to transform the feature planes. After training, feature planes are normalized and converted to 16-bit integers, corresponding to the YUV 16-bit format, and are encoded by a standard video codec.
@@
8.5 Compact3D: Compressing Gaussian Splat Radiance Field Models with Vector Quantization (Compact3D)
@@
8.6 CompGS: Efficient 3D Scene Representation via Compressed Gaussian Splatting CompGS
The paper CompGS [32] proposes a hybrid primitive structure with anchor primitives to predict the attributes of coupled primitives, resulting in compact residual representations. A rate-constrained optimization scheme further enhances compactness by jointly minimizing both rendering distortion and bit rate. The bit rate of both anchor and coupled primitives is modeled by entropy estimation.
@@ 8.11 Compressed 3D Gaussian Splatting for Accelerated Novel View Synthesis (Compressed3D)
The authors propose a compressed 3D Gaussian representation [43] consisting of three main steps: 1. sensitivity-aware clustering, where scene parameters are measured according to their contribution to the training images and encoded into compact codebooks via sensitivity-aware vector quantization; 2. quantization-aware fine-tuning, which recovers lost information by fine-tuning parameters at reduced bit-rates using quantization-aware training; and 3. entropy encoding, which exploits spatial coherence through entropy and run-length encoding by linearizing 3D Gaussians along a space-filling curve. Furthermore, a renderer for the compressed scenes utilizing GPU-based sorting and rasterization is proposed, enabling real-time novel view synthesis on low-end devices.
@@ 9.1 Mini-Splatting: Representing Scenes with a Constrained Number of Gaussians (Mini-Splatting)
Mini-Splatting [13] enhances Gaussian distribution through Blur Split, which refines Gaussians in blurred regions, and Depth Reinitialization, which repositions Gaussians based on newly generated depth points, calculated from the mid-point of ray intersections with Gaussian ellipsoids, thus avoiding artifacts from alpha blending. For simplification, Intersection Preserving retains Gaussians with the greatest visual impact, while Sampling maintains geometric integrity and rendering quality, reducing complexity.